# Friction terms in multi-fluid description of heavy-ion collisions

### v1

@@ IV Shear viscosity
The results of our simulations including shear viscosity are shown in Fig. 4. As expected, the charged hadron multiplicity is significantly increased and reproduces the data nicely. At the same time, the shear viscosity does not seem to affect the distribution of net protons too much. The main effect is a slight narrowing of the $dN_{p-\bar{p}}/dy$ distribution, which agrees with the idea of weaker longitudinal expansion. Overall, the inclusion of shear viscosity in the description reduces the tension between the reproduction of observed $dN_{\rm ch}/d\eta$ and $dN_{p-\bar{p}}/dy$.
Figure 5: Best fit to the experimental data, achieved with the friction terms settings $\alpha=0.8$, $\beta=0.1$, $\xi_{f\alpha}=0.1$, $\xi_{pt}=1.0$. For $\sqrt{s_{\rm NN}}=62.4$ GeV, an alternation $\alpha=0.75$, $\xi_{pt}=0.8$ is used. Fluids are evolved with temperature- and baryon chemical potential dependent shear viscosity from Eq. (18). Data are by the PHOBOS PHOBOS:2010eyu , NA49 NA49:1998gaz ; Blume:2007kw and BRAHMS BRAHMS:2009wlg collaborations.
@@ V Further data comparison
Up to this point, to gain insight in the roles of each of the model parameters in terms of physics and phenomenology, we concentrated on the Au+Au collisions at $\sqrt{s_{NN}}=19.6\,$GeV. Next, to test the predictive power of our model, we calculated the charged particle pseudorapidity distributions $dN_{\rm ch}/d\eta$, net-proton rapidity distributions $dN_{p-\bar{p}}/dy$, identified particle $p_{T}$-distributions ($\pi^{-}$, $K^{-}$ and protons) and $p_{T}$-differential elliptic flow $v_{2}(p_{T})$ of charged hadrons for the best parameter combinations at four different collision energies: $7.7$, $19.6$, $39$ and 62.462.4\,GeV.
The pseudorapidity and rapidity distributions are shown in Fig. 5. Data for the net-proton rapidity distribution is sparse, but for $7.7$ and 62.462.4\,GeV it has been measured by the NA49 NA49:1998gaz ; Blume:2007kw and BRAHMS BRAHMS:2009wlg collaborations. Additionally, data for the charged hadron distribution in pseudo-rapidity is available at 62.462.4\,GeV owing to PHOBOS PHOBOS:2010eyu . The plots show results for the parameter combination that best described experimental data. For 7.77.7\,GeV and 19.619.6\,GeV, these parameters are the same, and mentioned above, but for 62.462.4\,GeV $\alpha=0.75$ and $\xi_{\rm pt}=0.8$ led to better fit. We note that in previous works Ivanov:2013wha ; Cimerman:2023hjw using IMS friction, the parameters $\xi_{\rm pt}$ and $\xi_{\rm f}$ were modeled to depend on the CMS energy of the microscopic scattering processes, meaning they would change throughout the evolution. We use here a much weaker dependence only on the $\sqrt{s_{\rm NN}}$ of the entire nucleus-nucleus collision system. Our longitudinal distributions are in reasonable agreement with experimental data at all three collision energies. The model correctly reproduces the increase and widening of the charged hadron distribution with increasing energy. It also adequately reproduces in the net-proton distribution a transition from having almost no double peak structure at 7.77.7\,GeV to exhibiting two strongly separated peaks at 62.462.4\,GeV.
@@
Second, in Cimerman:2023hjw the calculated elliptic flow clearly overshot expectations, but inclusion of viscosity has brought it down to the vicinity of the data, as seen in the right plot of Fig. 6. We reproduce the data up to $p_{T}\approx 1$ GeV, but at larger values of $p_{T}$ our calculation suppresses $v_{2}$ too much, which again may indicate the need of bulk viscosity or improved $\eta(T,\mu)$. Nevertheless, it must be remembered that all the parameters of our model were fixed by comparison to rapidity and pseudorapidity distributions only, and our $p_{T}$ and elliptic flow results are predictions. From that point of view, our results are reasonable, and provide a good basis for further refinement of our model.
@@ VI Conclusion
We have performed scans of the four parameters of charge transfer friction to develop an understanding of their role in the physics at play. All parameters affect both particle production and baryon stopping via the amount of energy in the fireball, deceleration of target and projectile, expansion of the fireball, and/or the hadro-chemistry of the fluids. We did not find a parameter combination that would result in a simultaneous agreement with the experimental data for both the pseudo-rapidity density of all charged hadrons and the rapdidiy density of net protons. Cases that correctly reproduce the net-proton distribution would persistently underestimate charged hadron multiplicity. This leaves room for shear viscosity in the internal evolution of the fluids to increase entropy production. As we observed, inclusion of shear led to good agreement with data even across collision energies.
@@ Acknowledgements.
CW has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant number: 101089093 / project acronym: High-TheQ). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. IK acknowledges support by the Czech Science Foundation under project No. 25-16877S. PH was supported by the program Excellence Initiative–Research University of the University of Wrocław of the Ministry of Education and Science. Computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic.
@@ Appendix A Derivation of friction terms
The friction terms describing the transfer of energy and momentum between fluids can be derived from an underlying microscopic theory. We follow Satarov:1990uc ; Mishustin:1991sp ; Ivanov:2005yw in deriving them from kinetic theory, but we will make different modeling assumptions and point out where they play a role.
@@
$\partial_{\mu}N_{I}^{\mu}$ $=\int\mathrm{d}\Gamma_{I}~C_{I}$ (24)
@@
The collision kernels can now be expressed in terms of scattering cross-sections. Each component is considered to be made up of one type of particle, resulting in one specific cross-section. In the case that the fluids are mixtures, each fluid has to be considered to be made up of multiple components, but we will not model that here. We will also consider classical statistics, meaning that no quantum statistical factors occur in the collision kernel.
@@
$\sqrt{\left(p_{J,\mu}p_{K}^{\mu}\right)^{2}-m_{J}^{2}m_{K}^{2}}~\frac{\mathrm{d}\sigma_{JK\to IX}}{\mathrm{d}\Gamma_{I}}$
@@
$\sqrt{\left(p_{I,\mu}p_{J}^{\mu}\right)^{2}-m_{I}^{2}m_{J}^{2}}~\sigma_{IJ\to X}$
@@
In order to fix on top of baryon transfer also the energy balance of the target-projectile friction, technically we would need full information on $\mathrm{d}\sigma_{NN\to NX}/(\mathrm{d}y\mathrm{d}E)$, which is not available. Instead, we make the simplifying assumption that the part of the incoming energy $E_{0}$ that is not transferred to longitudinal momentum of the outgoing nucleon, $m_{N}\cosh(y)$, a fixed fraction will be transferred to the fireball, while the rest remains with the outgoing nucleon as transverse momentum. In other words,
@@
Furthermore, we assume for $\mathrm{d}\sigma_{NN\to NX}/\mathrm{d}y$ a functional form that can approximately describe the data but also allows to solve the occurring momentum integrals analytically.
$\frac{\mathrm{d}\sigma_{NN\to NX}}{\mathrm{d}y}\propto 1+6\frac{\cosh(10y)}{\cosh(10y_{p})}$ (31)
@@
$\frac{\mathrm{d}\sigma_{NN\to NX}}{dy}$ $=\frac{\sigma_{NN\to NX}}{2A}\left[1+6\frac{\cosh(10y)}{\cosh(10y_{p})}\right]$ (37)
@@
$+\frac{1}{2}(u_{p}^{\nu}+u_{t}^{\nu})\bar{\sigma}_{E}(s_{pt})+u^{\nu}_{p}\bar{\sigma}_{R}(s_{pt})\Big]\;.$ (53)
@@ Appendix B Fireball friction with thermal spread
$\tilde{f}_{f}=\frac{g_{\pi}}{\exp(\tilde{p}^{0}/T_{f})-1}\;,$ (57)
@@
and again analogously for $R_{B,p}$ without the factor of $\tilde{u}_{p}^{\nu}$ and dividing by $m_{N}$. The remaining integral has to be computed numerically. In the form of (59), it is obvious that this integral is a Lorentz scalar, so the transformation back to the CMS frame of the full collision system simply amounts to replacing $\tilde{u}_{p}^{\nu}$ by ${u}_{p}^{\nu}$.
@@ Appendix C Unification procedure
$\lambda=$ $\begin{cases}0\;,&v_{pf}>v_{T,p}\\ \exp\Big[\big(1-{v_{T,p}^{2}}/{v_{pf}^{2}}\big)^{-1}\Big]\;,&v_{pf}\leq v_{T,p}\end{cases}$ (66)

### v2

@@ IV Shear viscosity
The results of our simulations including shear viscosity are shown in Fig. 4. As expected, the charged hadron multiplicity is significantly increased and reproduces the data nicely. At the same time, the shear viscosity does not seem to affect the distribution of net protons too much. The main effect is a slight narrowing of the $dN_{p-\bar{p}}/dy$ distribution, which agrees with the idea of weaker longitudinal expansion. Overall, the inclusion of shear viscosity in the description reduces the tension between the reproduction of observed charged hadron pseudorapidity distribution $dN_{\rm ch}/d\eta$ and net-proton rapidity distribution $dN_{p-\bar{p}}/dy$.
Figure 5: Best fit to the experimental data at three different collision energies, achieved with the friction terms settings $\alpha=0.8$, $\beta=0.1$, $\xi_{f\alpha}=0.1$, $\xi_{pt}=1.0$. For $\sqrt{s_{\rm NN}}=62.4$ GeV, an alternation $\alpha=0.75$, $\xi_{pt}=0.8$ is used. Fluids are evolved with temperature- and baryon chemical potential dependent shear viscosity from Eq. (18). Data are by the PHOBOS PHOBOS:2010eyu , NA49 NA49:1998gaz ; Blume:2007kw and BRAHMS BRAHMS:2009wlg collaborations.
@@ V Further data comparison
Up to this point, to gain insight in the roles of each of the model parameters in terms of physics and phenomenology, we concentrated on the Au+Au collisions at $\sqrt{s_{NN}}=19.6\,$GeV. Next, to test the predictive power of our model, we calculated the charged hadron pseudorapidity distributions $dN_{\rm ch}/d\eta$, net-proton rapidity distributions $dN_{p-\bar{p}}/dy$, identified particle $p_{T}$-distributions ($\pi^{-}$, $K^{-}$ and protons) and $p_{T}$-differential elliptic flow $v_{2}(p_{T})$ of charged hadrons for the best parameter combinations at four different collision energies: $7.7$, $19.6$, $39$ and 62.462.4\,GeV.
The pseudorapidity and rapidity distributions are shown in Fig. 5. Data for the net-proton rapidity distribution is sparse, but for $7.7$ and 62.462.4\,GeV it has been measured by the NA49 NA49:1998gaz ; Blume:2007kw and BRAHMS BRAHMS:2009wlg collaborations. Additionally, data for the charged hadron distribution in pseudorapidity is available at 62.462.4\,GeV owing to PHOBOS PHOBOS:2010eyu . The plots show results for the parameter combination that best described experimental data. For 7.77.7\,GeV and 19.619.6\,GeV, these parameters are the same, and mentioned above, but for 62.462.4\,GeV $\alpha=0.75$ and $\xi_{\rm pt}=0.8$ led to better fit. We note that in previous works Ivanov:2013wha ; Cimerman:2023hjw using IMS friction, the parameters $\xi_{\rm pt}$ and $\xi_{\rm f}$ were modeled to depend on the CMS energy of the microscopic scattering processes, meaning they would change throughout the evolution. We use here a much weaker dependence only on the $\sqrt{s_{\rm NN}}$ of the entire nucleus-nucleus collision system. Our longitudinal distributions are in reasonable agreement with experimental data at all three collision energies. The model correctly reproduces the increase and widening of the charged hadron distribution with increasing energy. It also adequately reproduces in the net-proton distribution a transition from having almost no double peak structure at 7.77.7\,GeV to exhibiting two strongly separated peaks at 62.462.4\,GeV.
@@
Second, in Cimerman:2023hjw the calculated elliptic flow clearly overshot expectations, but inclusion of viscosity has brought it down to the vicinity of the data, as seen in the right plot of Fig. 6. We reproduce the data up to $p_{T}\approx 1$ GeV, but at larger values of $p_{T}$ our calculation suppresses $v_{2}$ too much, which again may indicate the need of bulk viscosity or improved shear viscosity parametrization $\eta(T,\mu)$. Nevertheless, it must be remembered that all the parameters of our model were fixed by comparison to rapidity and pseudorapidity distributions only, and our $p_{T}$ and elliptic flow results are predictions. From that point of view, our results are reasonable, and provide a good basis for further refinement of our model.
@@ VI Conclusion
We have performed scans of the four parameters of charge transfer friction to develop an understanding of their role in the physics at play. All parameters affect both particle production and baryon stopping via the amount of energy in the fireball, deceleration of target and projectile, expansion of the fireball, and/or the hadro-chemistry of the fluids. We did not find a parameter combination that would result in a simultaneous agreement with the experimental data for both the pseudorapidity density of all charged hadrons and the rapidity density of net protons. Cases that correctly reproduce the net-proton distribution would persistently underestimate charged hadron multiplicity. This leaves room for shear viscosity in the internal evolution of the fluids to increase entropy production. As we observed, inclusion of shear led to good agreement with data even across collision energies.
@@ Acknowledgements.
CW has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant number: 101089093 / project acronym: High-TheQ). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. IK acknowledges support by the Czech Science Foundation under project No. 25-16877S. PH was supported by the program Excellence Initiative–Research University of the University of Wrocław of the Ministry of Education and Science. Computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic, and by Wroclaw Centre for Networking and Supercomputing (http://wcss.pl).
Data availability
The data that support the findings of this article are openly available Zenodo .
@@ Appendix A Derivation of friction terms
The friction terms describing the transfer of energy and momentum between fluids can be derived from an underlying microscopic theory. We follow Refs. Satarov:1990uc ; Mishustin:1991sp ; Ivanov:2005yw in deriving them from kinetic theory, but we will make different modeling assumptions and point out where they play a role.
@@
$\partial_{\mu}N_{I}^{\mu}$ $=\int\mathrm{d}\Gamma_{I}~C_{I}.$ (24)
@@
The collision kernels can now be expressed in terms of scattering cross-sections. Each component is considered to be made up of one type of particle, resulting in one specific cross-section. In the case that the fluids are mixtures, each fluid has to be considered to be made up of multiple components, but we will not model that here. We will also consider classical statistics, meaning that no quantum statistical factors occur in the collision kernel:
@@
$\sqrt{\left(p_{J,\mu}p_{K}^{\mu}\right)^{2}-m_{J}^{2}m_{K}^{2}}~\frac{\mathrm{d}\sigma_{JK\to IX}}{\mathrm{d}\Gamma_{I}},$
@@
$\sqrt{\left(p_{I,\mu}p_{J}^{\mu}\right)^{2}-m_{I}^{2}m_{J}^{2}}~\sigma_{IJ\to X}.$
@@
In order to fix on top of baryon transfer also the energy balance of the target-projectile friction, technically we would need full information on $\mathrm{d}\sigma_{NN\to NX}/(\mathrm{d}y\mathrm{d}E)$, which is not available. Instead, we make the simplifying assumption that a fixed fraction of the part of the incoming energy $E_{0}$ that is not transferred to longitudinal momentum of the outgoing nucleon, $m_{N}\cosh(y)$, will be transferred to the fireball, while the rest remains with the outgoing nucleon as transverse momentum. In other words,
@@
Furthermore, we assume for $\mathrm{d}\sigma_{NN\to NX}/\mathrm{d}y$ a functional form that can approximately describe the data but also allows to solve the occurring momentum integrals analytically:
$\frac{\mathrm{d}\sigma_{NN\to NX}}{\mathrm{d}y}\propto 1+6\frac{\cosh(10y)}{\cosh(10y_{p})}.$ (31)
@@
$\frac{\mathrm{d}\sigma_{NN\to NX}}{dy}$ $=\frac{\sigma_{NN\to NX}}{2A}\left[1+6\frac{\cosh(10y)}{\cosh(10y_{p})}\right].$ (37)
@@
$+\frac{1}{2}(u_{p}^{\nu}+u_{t}^{\nu})\bar{\sigma}_{E}(s_{pt})+u^{\nu}_{p}\bar{\sigma}_{R}(s_{pt})\Big]\;.$
@@ Appendix B Fireball friction with thermal spread
$\tilde{f}_{f}=\frac{g_{\pi}}{\exp(\tilde{p}^{0}/T_{f})-1}\;.$ (56)
@@
and again analogously for $R_{B,p}$ without the factor of $\tilde{u}_{p}^{\nu}$ and dividing by $m_{N}$. The remaining integral has to be computed numerically. In the form of Eq. (58), it is obvious that this integral is a Lorentz scalar, so the transformation back to the CMS frame of the full collision system simply amounts to replacing $\tilde{u}_{p}^{\nu}$ by ${u}_{p}^{\nu}$.
@@ Appendix C Unification procedure
$\lambda=$ $\begin{cases}0\;,&v_{pf}>v_{T,p},\\ \exp\Big[\big(1-{v_{T,p}^{2}}/{v_{pf}^{2}}\big)^{-1}\Big]\;,&v_{pf}\leq v_{T,p},\end{cases}$ (65)